A voltmeter has a resistance G and range V. Calculate the resistance to be used to extend its range to nV.

Grade:12

2 Answers

Sher Mohammad IIT Delhi
9 years ago
The we require n resistance so , total resistance is nG
Sher Mohammad IIT Delhi
9 years ago
To extend the range of a voltmeter we add a resistance of more value in series with the meter.
value of added resistance(shunt) Rs=Rm+V'/ImHere V' is the new range, Rmis resistance of voltmeter and Imis maximum value of current through voltmeter that is Vm/Rm
for this case Rm=G Vm=V and V'=nV
so Rs=G+nG=(1+n)G

Sher Mohammad
faculty askiitians